# Tightrope-walking rationality in action: Feyerabendian insights for the foundations of quantum mechanics

### v2

@@ 3 Realism vs instrumentalism, in the interpretation of QM
Third, a realist methodology, i.e. the adoption of realism as a methodological principle has some further important methodological implications. In essence, one would say that it simply amounts to ”taking scientific theories seriously” for what they seem to say about the world, and explore their consequences in full, even when counterintuitive and conceptually weird, or even in contradiction with observations, being ready to reinterpret and modify the observational concepts if it seems useful to do so. Considering theories in isolation, this is it. When considering more than one theory at once, on the other hand, there is more. The natural consequence of methodological realism is that different theories (about the same physical domain), taken seriously and pushed to their extreme consequences, will in general conflict radically in their underlying (suggested) ontology. Insisting on a realist attitude, instead of retreating to an instrumentalist one (i.e. noticing the radical conflict in their core ingredients but being content with their agreement with observations, which of course only forces to accept a very limited part of them), implies adopting a further key methodological principle, at the core of PKF’s philosophy of science: proliferation (see the introduction to part 2 of [1] and chapters 3 and 4 in [35]. We need to be ready to pursue equally several competing theories, even when in radical conflict with each other and based on contradicting principles, coping with the contradiction and aiming for scientific progress and better understanding, rather than unity and agreement at the expenses of the potential progress, until the greater fruitfulness of one of them is established (and, PKF would argue, even afterwards). This proliferation of theories and research programmes, even contradicting one another at their core, is beneficial. There are many arguments for this conclusion, which was forcefully pushed by PKF, but has an earlier origin (dating back, as PKF himself notices, to J. Stuart Mill) and an independent academic lineage. Most of these arguments focus on the empirical assessment of theories and the discovery of observational anomalies, being often the direct result of the existence of alternatives to the given theory, and being at times invisible without such alternatives (see [36]). The importance of proliferation, however, can also be argued for in the context of non-empirical theory assessment [11], although of course on a different basis. None of these argument, in either context, is conclusive and the principle of proliferation remains controversial. See [37] for a review and more references.
@@
At this point, one could however feel that, once the arena of the debate between realism and instrumentalism has become that of scientific methodology, one could and probably should use the same arguments in favor of instrumentalism, rather than realism, as PKF does. In other words, PKF dismissal of ”instrumentalism” is correctly directed at dogmatic, positivist, excessively restrictive versions of it, but risks removing a useful methodological tool (contrary to his general stated goals). This would also follow naturally from PKF’s own methodological anarchism, in fact.
@@
Indeed, modern work in the foundations of QM has benefited greatly from the methodological advantages of ”instrumentalism”, in particular in its closely related version of ”operationalism”. The latter can be understood as the focus on the elements of QM that directly correspond to or follow from what agents (experimentalists or observers) actually do in the lab (intended in an idealized sense) when they manipulate (quantum) physical systems, somehow understanding the theory as a way to encode efficiently the results of such manipulations and its core principles as to be identified as stemming from them. This way of proceeding is most apparent in all the work on quantum foundations that proceeds in parallel with advances in quantum information and quantum computation, which provide at once a testing ground of foundational ideas and a source of inspiration for clarifying the foundations of QM. It can be seen in a lot of recent reconstructions of QM from operational principles, e.g. [38, 39] among many others, as well as in the development of a number of information-theoretic QM interpretations, e.g. [40, 41, 42] to mention just a few, which indeed are best understood as ”neo-Copenhagen”, inspired and further developed from the Bohr’s tradition [43]. We will say more about these modern developments in the next section. For now, it should be clear that ”instrumentalism” and ”operationalism” can be very fruitful methodological positions. Indeed, this is also how most of the scientists involved in the cited developments use them, without necessarily endorsing them as general philosophical positions, a convenient conceptual framework to force ourselves to focus on observable, testable, ”experienceable” aspects of scientific theories, and thus bring to the forefront some of their essential elements.
@@
In this context, PKF again contrasts the philosophers’ attitude to the physicists’ more pragmatic one, praising the latter, stressing again that physics arguments trump philosophical ones, however convincing. We thus come back to the discussion and conclusions of the previous section, on the general relation between philosophy (of physics) and physics. It is on this basis, as we have already noticed, that PKF came to gradually appreciate more and more Bohr’s philosophical and physical position, albeit an ”instrumentalist” one. And it is in the context of a reassessment of Bohr’s position that PKF offers further interesting insights on the foundations of QM, of a more specific kind (i.e. not having to do with the general discussion on realism vs instrumentalism, and the nature of physical theories). Some of them are by now standard understanding, accepted points in the domain of quantum foundations, and probably not because of a direct influence of PKF’s writings (whose impact on the philosophy of QM appears to us to have been rather limited). However, we find extremely interesting to notice their presence already in his (early) work. We now turn to a discussion of these insights.
@@ 4 Specific insights into the foundations of QM
The insights about QM from PKF that we find particularly interesting do not concern directly specific advances in the formalism or its use, even though in his writings one can find a number of careful analyses of QM results and some very useful comments about the quantum-classical transition, the relation between classical and quantum mechanics, and even some elements that seem to anticipate modern decoherence.
@@
The main one is a clear understanding of quantum states as relational in nature, rather than individually associated to physical systems as intrinsic (encoding of) properties. In QM the states associated to a physical system should always be understood in relation to a context, which in turn can be seen as corresponding to an observer or, more precisely, a measurement context. Because of this fundamentally relational nature, it becomes impossible to separate observer and observed system, and go back to an understanding of ’phenomena’ as revealed by observation but independent of the act of observation, or, more generally, the interaction between two (observer + observed) systems. This relational nature is not to be misunderstood (as often done in popular accounts of QM) as meaning that the interaction with the observer or the measurement apparatus, i.e. its embodiment, affects or modifies or disturbs the properties of the observed system, to which one associates quantum states. This picture assumes that the two systems in interaction and their properties can be characterized separately and their properties separately attributed. The message of QM, according to Bohr (mediated by PKF), is much stronger. Phenomena cannot be separated in individual components pertaining to the observed system and those pertaining to the observer (or measurement context). Using PKF’s words, in a section titled explicitly “The relational character of quantum states”:
[Complementarity] takes position, momentum, and all the other dynamical variables out of the individual physical system and attributes them to the experimental arrangement. The only properties which remain in the object and which can be attributed to it independently of the situation are mass, charge, baryon-number and similar ’non-dynamical’ characteristics. We shall say that complementarity asserts the relational character not only of probability, but of all dynamical magnitudes. One of the consequences of this procedure is that ’phenomena’ cannot be ’subdivided’ ([2], 50), which just means that dynamical magnitudes cannot be separated from the conditions of their application: ’this crucial point. . . implies the impossibility of any sharp separation between the behaviour of atomic objects and the interactions with the measuring instrument which serve to define the conditions under which the phenomena appear’ ([12], 39n°). Another consequence is that there arises an ’ambiguity in ascribing customary physical attributes to atomic objects’ ([12], 51) - for these attributes no longer apply to the object per se but to the whole experimental arrangement, different assertions being made in different contexts. Complementarity also offers a rough quantitative estimate of the extent to which experimental arrangements allowing for the application of certain notions (such as position) still permit the application of other notions (such as momentum). I am of course referring to the uncertainty principle which describes objective relations between states of systems in different experimental arrangements. - [1], page 260-261
@@
Now, it is clear that, as we discussed in the previous section, in Bohr we find a clear emphasis on the classical elements that we use to characterize the measurement context and thus the ”observing system” in ordinary applications of QM; and it is similarly clear that this emphasis brings Bohr’s position closer to instrumentalism. However, it is also clear that it is far from any strong or naive instrumentalism and that, while referring heavily to the results of measurements and observations, it does not take them conceptually for granted, as primitive notions not requiring a careful analysis, so much so that Bohr’s analysis leads indeed to a drastic revision, in QM, of what we mean by phenomena.
@@
This blurred distinction, in fact, does not save any naive realist position either, in the interpretation of QM put forward by Bohr (as understood by PKF). In this interpretation, what Bohr is doing is to propose a drastic revision of ”classical realism”, where the label ”classical” refers both to the picture of the world inherited by classical Newtonian mechanics and to traditional, strong forms of philosophical or scientific realism. It is not instrumentalism, but rather it should be understood as a starting point for developing a new form of realism.
@@
One relevant aspect is the following. It is thanks to this somewhat relaxed version of realism, following from a contextual, relational interpretation of quantum states, that Bohr’s view solves the EPR paradox without renouncing neither locality nor the completeness of QM. This transpires in full form when studying the Einstein-Bohr debate (e.g. starting from the Solvay conference [49]), but it hinges on this specific departure point from classical realism. In fact, this provides the key element for a general template for a solution of the EPR paradox or the interpretative conundrum offered by Bell’s theorem, which will then be followed and further articulated (in different ways) by modern QM interpretation in the ”neo-Copenhagen” tradition.
@@
Finally, one sees that Bohr’s suggestions provide an immediate solution to the famous case of Einstein, Podolsky and Rosen (EPR). [..] Einstein, Podolsky and Rosen assume without argument that what we determine when all disturbance has been eliminated is a property of the system examined that belongs to it irrespective of the conditions of the experiment (this is the content of their famous ’criterion of reality’). As opposed to this Bohr maintains, in accordance with his hypothesis, that quantum-mechanical state descriptions are relations between systems and experimental arrangement and are therefore dependent on the latter. It is easily seen how this assumption deals with EPR. For while a property cannot be changed except by direct interference with the system that possesses the property, a relation can be changed without such interference […] Hence, lack of physical interference excludes change of state only if it has already been established that positions and momenta and other magnitudes are properties of systems, belonging to them under all circumstances. But just this assumption had to be rejected and had to be replaced by Bohr’s hypothesis; and it had to be rejected because of difficulties which are quite independent of the case of EPR. Bohr’s solution of this case is therefore not only not ad hoc, but it adds further support to the idea of complementarity. - [1], page 292-293
@@
Next, the relational interpretation of quantum states, as we have said, injects a radical element of context-dependence in any realist perspective on physical theories [^9] We will discuss further this weakening of realism in the following. Here we notice that this is a central aspect, both profound and problematic, of several modern interpretations of quantum mechanics in various ways linked to Bohr’s and maintaining the relational character of quantum states. This is well recognized and subject of an extensive debate, at both physical and philosophical level, in the literature on quantum foundations. For one example among many, focusing on one specific interpretation, see [50].. This is a strong and rather obvious link between Bohr’s philosophy, and PKF’s reading of it, and modern developments in philosophy of science like structural realism [51], with its emphasis on relations as the ontological substratum of the world, but even more like the already cited perspectival realism of Giere [27, 28] and, more recently, Massimi [32]. The connections with the latter are probably stronger, also due to the lesser relevance of metaphysical considerations both in Bohr and modern perspectival realism (as already mentioned in the discussion of Giere’s methodological realism).
@@
It is also important to emphasize that, like in these modern versions of scientific realism, also in Bohr the relational nature of quantum states does not lead to any radical subjectivism or to the insertion of necessarily mind-related aspects in the description of nature. While the ”observer system” necessarily enters the characterization of quantum states and thus quantum phenomena always refer, at least implicitly, to a subject or agent, both these notions are entirely naturalised and objectified, in Bohr, and embodied in a measurement apparatus (described in classical terms). In particular, there is no role to play, in Bohr’s QM, for consciousness or higher-level agency, or for the personal experience of subjects. Even the notions of ”knowledge” and ”information” acquired by the observer in her interaction with the observed physical system (so relevant for some of the modern neo-Copenhagen interpretations) are not relevant in Bohr’s view. We are dealing, we could say, with a ”fully objectified subject” [^10] Here one can see maybe some traces of Whitehead’s thinking, already linked to the conceptual challenges of quantum mechanics in [52]. Along the same interpretation of PKF’s thinking, see also [53].. This should also be clear in light of the analogy, used by Bohr himself, and emphasized by PKF in the quote reported above, between the relative nature of quantum states, dependent on the measurement context, and the relative nature of mechanical quantities (e.g velocity) in relativistic physics, dependent on the reference frame. This is a parallel often made, by the way, by proponents of Relational Quantum Mechanics, and one could speculate that PKF could have seen, in this risky attempt at a middle ground between subjectivism and objectivism, in this weaker version of realism, another instance of tightrope-walking at a metaphysical and epistemological level.
@@
PKF is very careful in emphasizing this point, in defending Bohr’s views from Popper’s attack as excessively prone to subjectivism [^11] It is also interesting to note how PKF emphasizes the Kantian aspects of Bohr’s position and of the new relational understanding of quantum states and weaker realism, in line with what we also discussed above citing Hermann as well as recent scholars like Cuffaro and Crull.:
@@
We now come to the second point, Popper’s accusation of subjectivism. Bohr is interested in the interaction between the subject and the object and he also emphasizes the similarity between physics and psychology in which latter science ’we are continually reminded of the difficulty of distinguishing between subject and object’ ([8], 15 — original italics). But the ’subject’ in physics is for him not the consciousness of the observer but ’the agency’ used for observation, that is the material measuring instrument (including the body of the observer, and his sense organs); and the ’boundary’ dis- appears, in physics, not from between the consciousness of the observer and ’the world’; it disappears from between ’the [atomic] phenomena [and] the [material] agencies of observation’ ([8], 54). […] - [1], page 280-282
@@
So, the subject/observer in QM is fully embodied in a measurement context, but at the same time the object/observed looses necessarily its absolute, independent characterization.
@@
It remains for us to connect what has been said so far with the hypothesis of the relational character of the quantum-mechanical states which was formulated in section 4 above. First of all it should now be clear that this hypothesis is a physical hypothesis and not the consequence of some inarticulate philosophical faith. It is to be admitted that Bohr also had philosophical reasons to expect a hypothesis of this kind to be true. But we have seen that this philosophy did not prevent him from exploring alternatives; and we have also seen that it was the refutation of these alternatives taken together with difficulties such as those mentioned in section 5 which led him back to his original ideas and convinced him, almost to his surprise, of the correctness of his more philosophical outlook. Secondly, the hypothesis does not introduce any subjective element but concerns the physical situation only […]. The hypothesis has certain similarities with Kantianism, assuming that nature (and the word is here used in almost the same sense as in Kant) depends on our categories and forms of perception. The difference is that the physical interactions involved in any act of cognition are taken into account and that consequences for epistemology are drawn from such physical considerations […]. - [1], page 280-282
Thus, this is not the result of a philosophical analysis; it is the conclusion to which Bohr is led by physical arguments and results, guided by a philosophical stimulus but after having explored seriously the alternatives.
@@
Finally, it is evident that the ’observer’ and his ’knowledge’ nowhere enter the scene. All that is asserted is that there are objective conditions which do not allow for the application of certain magnitudes (and of the theories relating these magnitudes to each other) irrespective of whether these objective conditions are now used for improving our knowledge or not. In order to make this feature of the hypothesis as clear as possible I shall now illustrate it, as did Bohr ([9] and [88]), with the help of the very similar situation in relativity. […] An object may be heavy in one frame, light in another, i.e. $M[0,R^{\prime}]$ may be $M[0,R"]$ and the value changes from one to the other depending on whether we consider the one or the other reference system. […] But again, it would be quite mistaken to try explaining the change in a ’mechanistic’ fashion, as the result of an interaction between 0 and R (as is attempted in the electron theory of H. A. Lorentz); and we must again recall Bohr’s warning not to confuse ’a mechanical disturbance of the system under consideration’ with an ’influence on the very conditions’ of the experiment. - [1], page 280-282
And all of the above, thus QM, leads to a novel form of relational objectivity.
@@
It is also in this ”relational objectivity”, that one can see most clearly the connection between the relative nature of quantum states proposed by Bohr (that is, ”complementarity”) and the proposal made by Popper (and others) of an objective nature of probabilities (including quantum probabilities) in terms of propensities.
@@
Before we expand on this connection, let us stress that it is inevitable to inquiry about the connection between how one interprets QM states and how one interprets probabilities, and a coherence between the two can even be demanded as a consistency condition of one’s overall picture of the world. On the one hand, all that QM predicts are probabilities of results of future measurements or, in a more ”realist”  take, the probabilities that the quantum systems (will be shown to) possess specific properties; thus also how we understand probabilities immediately enters how we interpret QM and viceversa, generally speaking. On the other hand, more specifically, quantum states can be directly traded for a list of probability assignments, both mathematically (one can reformulate the whole of QM replacing quantum states with probability assignments for the possible values of the various observables characterizing the system) and conceptually (e.g. holding that quantum states are just computational tools for real but probabilistic properties of systems, or that they simply encode observer’s knowledge or information or belief about properties of systems, and it is this knowledge or information or belief that is real; that is, both in an objectivist and subjectivist understanding of probabilities). Thus, what interpretation of probabilities can be put in correspondence with Bohr’s views and ”orthodox” QM or with other QM interpretations? Viceversa, does a particular view of the nature of probabilities imply or favour a specific interpretation of QM?
@@
These are questions as difficult as they are interesting. Here we simply identify what PKF had to say about one specific instance of such relation, once more in a virtual debate with Popper, which apart from anything else shows how both Popper and PKF had already clear the strict relation between (the interpretation of) quantum states and probabilities.

### v3

@@ 3 Realism vs instrumentalism, in the interpretation of QM
Third, a realist methodology, i.e. the adoption of realism has some further important methodological implications. In essence, it seems to amount to ”taking scientific theories seriously” for what they seem to say about the world, and explore their consequences fully, even when counterintuitive and conceptually weird or even in contradiction with observations while being ready to reinterpret and modify the observational concepts if it seems useful to do so. When considering more than one theory at once, on the other hand, there is more to gain from a realist methodology. The natural consequence of methodological realism is that different theories (about the same physical domain), taken seriously and pushed to their extreme consequences, will usually conflict radically in their underlying (suggested) ontology. Insisting on a realist attitude, instead of retreating to an instrumentalist one (i.e. noticing the radical conflict in their core ingredients but being content with their agreement with observations, which of course only forces to accept a very limited part of them), implies adopting a further key methodological principle, at the core of PKF’s philosophy of science: proliferation (see the introduction to part 2 of [1], chapters 3 and 4 in [49], and [7]. We need to be ready to pursue equally several competing theories, even when in radical conflict with each other and based on contradicting principles, coping with the contradiction and aiming for scientific progress and better understanding, rather than unity and agreement at the expenses of the potential progress, until the greater fruitfulness of one of them is established (and, PKF would argue, even afterwards). This proliferation of theories and research programmes, even contradicting one another at their core, is beneficial. There are many arguments for this conclusion, which was forcefully pushed by PKF, but has an earlier origin (dating back, as PKF himself notices, to J. Stuart Mill [50]) and an independent academic lineage. Most of these arguments focus on the empirical assessment of theories and the discovery of observational anomalies, being often the direct result of the existence of alternatives to the given theory, and being at times invisible without such alternatives (see [51]). The importance of proliferation, however, can also be argued for in the context of non-empirical theory assessment [21], although of course on a different basis. None of these argument, in either context, is conclusive and the principle of proliferation remains controversial. See [52] for a review and more references.
@@
At this point, once the arena of the debate between realism and instrumentalism has become that of scientific methodology, one could and probably should use the same arguments in favor of instrumentalism, rather than realism, as PKF does. In other words, PKF dismissal of ”instrumentalism” is correctly directed at dogmatic, positivist, excessively restrictive versions of it, but risks removing a useful methodological tool (contrary to his general stated goals). This would also follow naturally from PKF’s own methodological anarchism, in fact.
@@
Indeed, modern work in the foundations of QM has benefited greatly from the methodological advantages of ”instrumentalism”, in particular in its closely related version of ”operationalism”. The latter can be understood as the focus on the elements of QM that directly correspond to or follow from what agents (experimentalists or observers) actually do in the lab (intended in an idealized sense) when they manipulate (quantum) physical systems, somehow understanding the theory as a way to encode efficiently the results of such manipulations and its core principles as to be identified as stemming from them. This way of proceeding is most apparent in all the work on quantum foundations that proceeds in parallel with advances in quantum information and quantum computation, which provide at once a testing ground of foundational ideas and a source of inspiration for clarifying the foundations of QM. It can be seen in a lot of recent reconstructions of QM from operational principles, e.g. [53, 54] among many others, as well as in the development of a number of information-theoretic QM interpretations, e.g. [55, 56, 57] to mention just a few, which indeed are best understood as ”neo-Copenhagen”, inspired and further developed from the Bohr’s tradition [58]. We will say more about these modern developments in the next section. For now, it should be clear that ”instrumentalism” and ”operationalism” can be very fruitful methodological positions. Indeed, this is also how most of the scientists involved in the cited developments use them, without necessarily endorsing them as general philosophical positions but as a convenient conceptual framework to force ourselves to focus on observable, testable, ”experienceable” aspects of scientific theories, and thus bring to the forefront some of their essential elements.
@@
In this context, PKF again contrasts the philosophers’ attitude to the physicists’ more pragmatic one, praising the latter, stressing again that physics arguments trump philosophical ones, however convincing. We thus come back to the discussion and conclusions of the previous section, on the general relation between philosophy (of physics) and physics. It is on this basis, as we have already noticed, that PKF came to gradually appreciate more and more Bohr’s philosophical and physical position, albeit an ”instrumentalist” one. And it is in the context of a reassessment of Bohr’s position that PKF offers further interesting insights on the foundations of QM, of a more specific kind (i.e. not having to do with the general discussion on realism vs instrumentalism, and the nature of physical theories). Some of them are by now standard understanding, accepted points in the domain of quantum foundations, and probably not because of a direct influence of PKF’s writings (whose impact on the philosophy of QM appears to have been rather limited). However, we find extremely interesting to notice their presence already in his (early) work. We now turn to a discussion of these insights.
@@ 4 Specific insights into the foundations of QM
The insights about QM from PKF that we find particularly interesting do not directly concern specific advances in the formalism or its use, even though in his writings one can find a number of careful analyses of QM results and some very useful comments about the quantum-classical transition, the relation between classical and quantum mechanics, and even some elements that seem to anticipate modern decoherence.
@@
The main one is a clear understanding of quantum states as relational in nature, rather than individually associated to physical systems as intrinsic (encoding of) properties. In QM, the states associated to a physical system should always be understood in relation to a context, which in turn can be seen as corresponding to an observer or, more precisely, a measurement context. Because of this fundamentally relational nature, it becomes impossible to separate observer and observed system, and go back to an understanding of ’phenomena’ as revealed by observation but independent of the act of observation, or, more generally, the interaction between two (observer + observed) systems. This relational nature is not to be misunderstood (as often done in popular accounts of QM) as meaning that the interaction with the observer or the measurement apparatus, i.e. its embodiment, affects or modifies or disturbs the properties of the observed system, to which one associates quantum states. This picture assumes that the two systems in interaction and their properties can be characterized separately and their properties separately attributed. The message of QM, according to Bohr (mediated by PKF), is much stronger. Phenomena cannot be separated in individual components pertaining to the observed system and those pertaining to the observer (or measurement context). Using PKF’s words, in a section titled explicitly “The relational character of quantum states”:
[Complementarity] takes position, momentum, and all the other dynamical variables out of the individual physical system and attributes them to the experimental arrangement. The only properties which remain in the object and which can be attributed to it independently of the situation are mass, charge, baryon-number and similar ’non-dynamical’ characteristics. We shall say that complementarity asserts the relational character not only of probability, but of all dynamical magnitudes. One of the consequences of this procedure is that ’phenomena’ cannot be ’subdivided’ ([2], 50), which just means that dynamical magnitudes cannot be separated from the conditions of their application: ’this crucial point. . . implies the impossibility of any sharp separation between the behaviour of atomic objects and the interactions with the measuring instrument which serve to define the conditions under which the phenomena appear’ ([12], 39n°). Another consequence is that there arises an ’ambiguity in ascribing customary physical attributes to atomic objects’ ([12], 51) - for these attributes no longer apply to the object per se but to the whole experimental arrangement, different assertions being made in different contexts. - [1], page 260-261
@@
Now, it is clear that, as we discussed in the previous section, in Bohr we find a clear emphasis on the classical elements that we use to characterize the measurement context and thus the ”observing system” in ordinary applications of QM. It is similarly clear that this emphasis brings Bohr’s position closer to instrumentalism. However, it is also clear that it is far from any strong or naive instrumentalism and that, while referring heavily to the results of measurements and observations, it does not take them for granted as primitive notions not requiring a careful analysis, so much so that Bohr’s analysis leads indeed to a drastic revision, in QM, of what we mean by ‘phenomena’.
@@
This blurred distinction, in fact, does not save any naive realist position either, in the interpretation of QM put forward by Bohr (as understood by PKF). In this interpretation, what Bohr is doing is proposing a drastic revision of “classical realism”, where the label ”classical” refers both to the picture of the world inherited by classical Newtonian mechanics and to traditional, strong forms of philosophical or scientific realism. It is not instrumentalism, but rather it should be understood as a starting point for developing a new form of realism.
@@
One relevant aspect is the following. It is thanks to this somewhat relaxed version of realism, following from a contextual, relational interpretation of quantum states, that Bohr’s view solves the EPR paradox without renouncing locality nor the completeness of QM. This transpires in full form when studying the Einstein-Bohr debate (e.g. starting from the Solvay conference [64]), but it hinges on this specific departure from classical realism. In fact, this provides the key element for a general template for a solution of the EPR paradox or the interpretative conundrum offered by Bell’s theorem, which will then be followed and further articulated (in different ways) by modern QM interpretation in the ”neo-Copenhagen” tradition.
@@
[..] Einstein, Podolsky and Rosen assume without argument that what we determine when all disturbance has been eliminated is a property of the system examined that belongs to it irrespective of the conditions of the experiment (this is the content of their famous ’criterion of reality’). As opposed to this Bohr maintains, in accordance with his hypothesis, that quantum-mechanical state descriptions are relations between systems and experimental arrangement and are therefore dependent on the latter. It is easily seen how this assumption deals with EPR. For while a property cannot be changed except by direct interference with the system that possesses the property, a relation can be changed without such interference […] Hence, lack of physical interference excludes change of state only if it has already been established that positions and momenta and other magnitudes are properties of systems, belonging to them under all circumstances. But just this assumption had to be rejected and had to be replaced by Bohr’s hypothesis; and it had to be rejected because of difficulties which are quite independent of the case of EPR. Bohr’s solution of this case is therefore not only not ad hoc, but it adds further support to the idea of complementarity. - [1], page 292-293
@@
Next, the relational interpretation of quantum states, as we have said, injects a radical element of context-dependence in any realist perspective on physical theories [^8] We will discuss further this weakening of realism in the rest of this section. Here we notice that this is a central aspect, both profound and problematic, of several modern interpretations of quantum mechanics in various ways linked to Bohr’s and maintaining the relational character of quantum states. This is well recognized and subject of an extensive debate, at physical and philosophical levels in the literature on quantum foundations. For one example among many, focusing on one specific interpretation, see [65].. This is a strong and rather obvious link between Bohr’s philosophy, and PKF’s reading of it, and modern developments in philosophy of science like structural realism [66], with its emphasis on relations as the ontological substratum of the world, but even more like the already cited perspectival realism of Giere [41, 42] and, more recently, Massimi [46]. The connections with the latter are probably stronger, also due to the lesser relevance of metaphysical considerations both in Bohr and modern perspectival realism (as already mentioned in the discussion of Giere’s methodological realism).
@@
It is also important to emphasize that, like in these modern versions of scientific realism, the relational nature of quantum states does not lead to any radical subjectivism or to the insertion of necessarily mind-related aspects in the description of nature, in Bohr’s perspective. While the ”observer system” necessarily enters the characterization of quantum states and thus quantum phenomena always refer, at least implicitly, to a subject or agent, both these notions are entirely naturalised and objectified, and embodied in a measurement apparatus (described in classical terms). In particular, there is no role to play for consciousness or higher-level agency, or for the personal experience of subjects. Even the notions of ”knowledge” and ”information” acquired by the observer in her interaction with the observed physical system (so relevant for some of the modern neo-Copenhagen interpretations) are not relevant in Bohr’s view. We are dealing, we could say, with a ”fully objectified subject” [^9] Here one can see maybe some traces of Whitehead’s thinking, already linked to the conceptual challenges of quantum mechanics in [67]. Along the same interpretation of PKF’s thinking, see also [68].. This should also be clear in light of the analogy, used by Bohr himself, and emphasized by PKF in the quote reported above, between the relative nature of quantum states, dependent on the measurement context, and the relative nature of mechanical quantities (e.g velocity) in relativistic physics, dependent on the reference frame. This is a parallel often made by proponents of Relational Quantum Mechanics, and one could speculate that PKF could have seen, in this risky attempt at a middle ground between subjectivism and objectivism, another instance of tightrope-walking at a metaphysical and epistemological level.
@@
PKF is very careful in emphasizing this point, in defending Bohr’s views from Popper’s attack as excessively prone to subjectivism [^10] It is also interesting to note how PKF emphasizes the Kantian aspects of Bohr’s position and of the new relational understanding of quantum states and weaker realism, in line with what we also discussed above citing Hermann as well as recent scholars like Crull and Cuffaro.:
@@
Bohr is interested in the interaction between the subject and the object and he also emphasizes the similarity between physics and psychology in which latter science ’we are continually reminded of the difficulty of distinguishing between subject and object’ ([8], 15 — original italics). But the ’subject’ in physics is for him not the consciousness of the observer but ’the agency’ used for observation, that is the material measuring instrument (including the body of the observer, and his sense organs); and the ’boundary’ dis- appears, in physics, not from between the consciousness of the observer and ’the world’; it disappears from between ’the [atomic] phenomena [and] the [material] agencies of observation’ ([8], 54). […] - [1], page 280-282
@@
So, the subject/observer in QM is fully embodied in a measurement context, but at the same time the object/observed necessarily looses its absolute, independent characterization. The absence of subjective (let alone mental/conscious) and epistemic elements, in this relational view, is repeatedly stressed by PKF:
@@
[the relational character of the quantum-mechanical states] does not introduce any subjective element but concerns the physical situation only […]. The hypothesis has certain similarities with Kantianism, assuming that nature (and the word is here used in almost the same sense as in Kant) depends on our categories and forms of perception. The difference is that the physical interactions involved in any act of cognition are taken into account and that consequences for epistemology are drawn from such physical considerations […]. - [1], page 280-282
and again:
@@
Finally, it is evident that the ’observer’ and his ’knowledge’ nowhere enter the scene. All that is asserted is that there are objective conditions which do not allow for the application of certain magnitudes (and of the theories relating these magnitudes to each other) irrespective of whether these objective conditions are now used for improving our knowledge or not. to make this feature of the hypothesis as clear as possible I shall now illustrate it, as did Bohr ([9] and [88]), with the help of the very similar situation in relativity. […] An object may be heavy in one frame, light in another, i.e. $M[0,R^{\prime}]$ may be $M[0,R"]$ and the value changes from one to the other depending on whether we consider the one or the other reference system. […] But again, it would be quite mistaken to try explaining the change in a ’mechanistic’ fashion, as the result of an interaction between 0 and R (as is attempted in the electron theory of H. A. Lorentz); and we must again recall Bohr’s warning not to confuse ’a mechanical disturbance of the system under consideration’ with an ’influence on the very conditions’ of the experiment. - [1], page 280-282
All of the above suggests a novel form of relational objectivity.
@@
It is also in this ”relational objectivity” that one can see the connection between the relative nature of quantum states proposed by Bohr (that is, ”complementarity”) and the proposal made by Popper (and others) of an objective nature of probabilities (including quantum probabilities) in terms of propensities, most clearly.
@@
Before we expand on this connection, let us stress that it is inevitable to inquire about the connection between how one interprets QM states and how one interprets probabilities, and a coherence between the two can even be demanded as a consistency condition of one’s overall picture of the world. On the one hand, all that QM predicts are probabilities of the results of future measurements or, in a more ”realist”  take, the probabilities that the quantum systems (will be shown to) possess specific properties. This also impacts how we understand probabilities immediately enters how we interpret QM and vice versa, generally speaking. On the other hand, more specifically, quantum states can be directly traded for a list of probability assignments, both mathematically (one can reformulate the whole of QM replacing quantum states with probability assignments for the possible values of the various observables characterizing the system) and conceptually (e.g. holding that quantum states are just computational tools for real but probabilistic properties of systems, or that they simply encode observer’s knowledge or information or belief about properties of systems, and it is this knowledge or information or belief that is real; that is, both in an objectivist and subjectivist understanding of probabilities). Thus, what interpretation of probabilities can be put in correspondence with Bohr’s views and ”orthodox” QM or with other QM interpretations? Vice versa, does a particular view of the nature of probabilities imply or favour a specific interpretation of QM?
@@
These questions are as difficult as they are interesting. Here we simply identify what PKF had to say about one specific instance of such relation, once more in a virtual debate with Popper, which apart from anything else shows how both Popper and PKF had already clear the strict relation between (the interpretation of) quantum states and probabilities.